HUBBLE SPACE TELESCOPE DAILY REPORT #2241 PERIOD COVERED: 0000Z (UTC) 10/30/98 - 0000Z (UTC) 11/02/98 Daily Status Report as of 306/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Thirteen Sets of WF/PC-2 7713 (WF/P-2 Cycle 7 Supplemental Darks pt3/3) The WF/PC-2 was used to make supplemental dark calibration images. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were completed as planned, and no anomalies were noted. 1.2 Completed Five Sets of STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed Twenty-Two Sets of NIC/1/2/3 7963 (Darks: Warming Up) The Near Infrared Camera (Cameras One, Two and Three) was used to monitor the detector dark current during the interval between the end of science observations and the exhaustion of cryogen and subsequent warming of the dewar to > 100K. The observations were completed as planned, and no anomalies were noted. 1.4 Completed Two Sets of NIC/2 7894 (Completing A Near-Infrared Search for Very Low Mass Companions to Stars within 10 pc) The Near Infrared Camera (Camera Two) was used to make observations searching for very low mass (VLM) companions to the stars CD-44D3045 and LHS271. This search will be the largest, most sensitive, volume-limited search for VLM companions ever undertaken. The observations were completed as planned, and no anomalies were reported. 1.5 Completed FOC/96 7685 (FOC F/96 UV Throughput Monitoring) The Faint Object Camera (f/96) was used to make observations of standard star BPM16274 in order to perform ultraviolet throughput monitoring. These observations are taken four time during the cycle. An internal calibration flat image was also taken. The observations were completed as scheduled, and no problems were reported. 1.6 Completed Thirteen Sets of STIS/CCD 8084 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation II) The Space Telescope Imaging Spectrograph (CCD) was used to make non-scripted parallel high galactic latitude observations as part of POMS test proposal. The observations were executed as scheduled, and no anomalies were noted. 1.7 Completed Six Sets of WF/PC-2 7909 (POMS Test Proposal: WF II Parallel Archive Proposal) The WF/PC-2 was used to make parallel observations as part of the POMS test proposal. This test proposal is designed to simulate scientific plans. The observations were completed as planned, and no problems were reported. 1.8 Completed Five Sets of NIC/2/3 8082 (NICMOS Pure Parallel Fall 98) The Near Infrared Camera (Cameras Two and Three) was used to make pure parallel observations while another instrument is prime. The observations were executed as scheduled, and no anomalies were noted. 1.9 Completed STIS/CCD/MA1/MA2 7939 (Determination of OI Emissions from Ganymede) The Space Telescope Imaging Spectrograph (CCD/MA1/MA2) was used to spectroscopically image the OI emissions from Ganymede. The Galileo mission has shown that Ganymede possesses an internal magnetic field, producing a local magnetosphere within the magnetosphere of Jupiter. It is highly desirable that these observations be obtained in calendar 1998 in order to best support the extended Galileo mission. The observations completed as scheduled with no reported problems. 1.10 Completed Two Sets of NIC/2 7227 (A Search for Low Mass/Sub-Luminous Companions to M-Stars) The Near Infrared Camera (Camera Two) was used to make a search for low mass/sub-luminous companions to low mass M class stars. Knowledge of stellar and sub-stellar masses and luminosities at and below the ~0.08 M{sun} H-burning limit is of fundamental importance. A dark calibration image was also taken. The observations were completed as planned, and no problems were reported. 1.11 Completed Five Sets of NIC/2 7226 (Search for Massive Jupiters) The Near Infrared Camera (Camera Two) was used to make a corographic search for massive planets around nearby, young main sequence stars. Because of the extreme youth of these objects, any low-mass brown dwarf and planetary companions will still be in a higher luminosity phase and thus easily detectable. A dark calibration image was also taken. The observations were executed as scheduled, and no anomalies were noted. 1.12 Completed WF/PC-2 6597 (High Resolution Imagery of Bipolar, Rotating, Episodic Jets) The WF/PC-2 was used to search for evidence of the existence of bipolar, rotating, episodic jets. These jets are characterized by symmetric, diametrically opposite, collimated flows with observable, radiative, bow-shocks around their leading edges. It is the spatial distributions of the resultant knots of ionized gas which indicate rotations of the bi-polar, episodic jets. The observations were completed as planned, and no anomalies were reported. 1.13 Completed Two Sets of NIC/2 7233 (Dust Disks Around Main Sequence Stars) The Near Infrared Camera (Camera Two) was used to make observations of the dust disk around main sequence stars. The observations were made with the coronagraph to minimize the effects of glare from the bright central star. A dark calibration image was also taken. The observations were completed as scheduled, and no anomalies were reported. 1.14 Completed STIS/CCD 7361 (The Nature of Nuclear Activity in Nearby Galaxies) The Space Telescope Imaging Spectrograph (CCD) was used to make observations of the nuclear region of galaxy N3368. The high resolution spectra will be used to search for broad components in H-Alpha and for velocity gradients, indicative of gas rotation or other ordered motions. The observations were executed as scheduled, and no anomalies were noted. 1.15 Completed WF/PC-2 7277 (Calibration of the Cepheid P-L Relation with Observations of the Maser-Host Galaxy NGC) The WF/PC-2 was used to make observations of the maser-host galaxy NGC-4258 to study its Cepheid variable stars. These data will help refine the calibration of the Cepheid period-luminosity relationship. The observations were completed as scheduled, and no problems were reported. 1.16 Completed Two Sets of NIC/2 7897 (Probing Planetary Formation Around Main-Sequence Stars: A Snapshot Survey) The Near Infrared Camera (Camera Two) was used to make snapshort survey images of the main sequence stars HR241 and HR1686 probing for planetary formation. The IR excess main-sequence stars host cool circumstellar dust disks, which are most likely replenished by the debris of unseen planetesimals. The observations were completed as scheduled, and no problems were noted. 1.17 Completed WF/PC-2 6734 (Revelaing the Stellar Populations of Andromeda IV) The WF/PC-2 was used to observe Andromeda IV, an enigmatic object, first identified in 1972 by proposing deep WFPC2 V and I imaging of And IV to V>~ 27, which will resolve stars down to M_V ~ +3, and determine main sequence turnoffs of <~ 5 * 10^9yr. The observations completed as planned. 1.18 Completed Four Sets of STIS/CCD 7912 (STIS Parallel Archive Proposal - Nearby Galaxies - Imaging and Spectroscopy) The Space Telescope Imaging Spectrograph (CCD) was used to make parallel observations of nearby galaxies. This survey will be useful to study the star formation histories, chemical evolution, and distances to these galaxies. These data will be placed immediately into the Hubble Data Archive. The observations were completed as with no problems. 1.19 Completed Three Sets of NIC/2/3 8063 (NICMOS Pure Parallel Sponge - Backup) The Near Infrared Camera (Cameras Two and Three) was used to make observations taken in parallel with the prime science observations. This proposal is PRE-the new FOM default. The observations were executed as scheduled, and no anomalies were noted. 1.20 Completed STIS/CCD 7333 (STIS Images of 48 Damped Ly-Alpha Galaxies) The Space Telescope Imaging Spectrograph (CCD) was used to make observations of the damped Lyman-Alpha galaxy Q0049+2820. This study will unify absorption-line studies, and faint galaxy imaging studies, thereby allowing a detailed treatment of the history of star formation, gas consumption, chemical enrichment, and assembly in the population of galaxies. The observations were completed as planned, and no problems were reported. 1.21 Completed NIC/2 7890 (The Youngest Stars: Circumstellar Structure, Binary and Origin of Jets) 4| The Near Infrared Camera (Camera Two) was used to make observations of the young star HH300-VLA which has circumstellar structure, and jets. The purpose is to obtain broadband 2.2 Mum images of the compact cavities illuminated by Herbig-Haro (HH) energy sources, and to obtain narrowband H-2 images to study how the HH jets originate at their sources. The observations were completed as planned, and no problems were reported. 1.22 Completed NIC/3 7892 (Narrow-band Imaging of BAL QSOs) The Near Infrared Camera (Camera Three) was used to make narrow band observations of the BL-Lac (BAL) QSO UN253. The detection of extended emission-line gas structures in BAL QSOs would offer another means of investigating the BAL QSO phenomenon as well as establishing a new probe of material in galaxies at high redshift. The observations were completed as planned, and no problems were reported. 1.23 Completed WF/PC-2 7387 (T-Tauri Star Snapshot Survey) The WF/PC-2 was used to make observations of the star HP-TAUG2 as part of a T-Tauri class star survey. The observations were executed as scheduled with no anomalies. 1.24 Completed WF/PC-2 6803 (Disk-Resolved Spectrophotometry of the Dark Side of Iapetus) The WF/PC-2 was used to obtain the first spectrum of pure dark material on Iapetus' Cassini Regio from near UV to near IR {335-1020 nm}, allowing a calibration of higher resolution, earth-based spectra of the leading side of Iapetus. The proposal was executed as planned with no problems. 1.25 Completed STIS/CCD/MA1/MA2 7269 (The 1997/98 Eclipse of VV Cephei) The Space Telescope Imaging Spectrograph (CCD, MA1 and MA2) was used to make ultraviolet spectrographic observations of the long-period eclipsing binary star HD208816 (a.k.a., VV-Cephei). This eclipse will provide an unprecedented opportunity to obtain spatially resolved information about the density, ionization and velocity structure in the outer atmosphere of the M supergiant primary. A wave length calibration observation was also taken. The observations were completed as scheduled, and no anomalies were noted. 1.26 Completed WF/PC-2 8061 (POMS Test Proposal: WF-II Backup Parallel Archive Proposal) This is a POMS test proposal designed to simulate scientific plans. The proposal was executed as planned, and no anomalies were reported. 1.27 Completed WF/PC-2 6677 (The Origin of the X-Ray Background: What are the Faintest X-Ray Galaxies?) The WF/PC-2 was used to make observations of the galaxy GAL-133525 which is from the deepest optically identified (ROSAT) X-ray survey yet. These images were taken to determine the morphologies of the remaining systems which we cannot study from the ground to determine whether the remaining non-QSO XRB might be explained by emission from interacting or disturbed spirals. The observations were completed as planned, and no anomalies were reported. 2.0 FLIGHT OPERATIONS SUMMARY 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 34 Successful : 34 Scheduled Reacquisitions : 15 Successful : 15 2.2 FHST Updates: Scheduled : 58 Successful : 58 2.3 Operations Notes: Using ROP SR-1, the SSR EDAC error counter was cleared eight times. PDB 53.C was made prime at 304/1720Z. No problems were encountered either before (when it operated in the sshadow mode, or after the transition. As documented in HSTAR 6748, the STIS MCE-1 reset at 303/171503Z while the vehicle was outside the SAA and while the low voltage was on. ROP NS-12 was executed to reset the STIS error counter. The MAMA-1 was recovered successfully via the SMS at the time of the next high voltage on commanding, or 305/073634Z. A NICMOS single bit error occurred at 305/062456Z. HSTAR 6750 was written. Per an operations request, the NICMOS internal log was dumped and the ground system limit was raised by one count at 305/0740Z. As documented in HSTAR 6751, the STIS MCE-2 reset at 305/162932Z while the vehicle was in an SAA interval and while the low voltage was on. ROP NS-12 was executed to reset the STIS error counter. The MAMA-2 was recovered successfully via the SMS at the time of the next high voltage on commanding. HSTAR 6754 describes a large roll error of 56.6 arc-sec that occurred at 306/062148Z during dual FGS guidance. A proposal to be described in the next report, NICMOS 7635, may have been affected. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH